In review for an exam, I've back-tracked to see where I got lost and here it is: with the ordered pair (0,5) and (4,0), the origin lies on both the x and y axis, its coordinates are (0,0). The origin is often labeled with the letter "o". How do you know the origin and where did the (0,0)come from? I hope this doesn't sound too stupid! Thanks in advance
Answer by Nate(3500)

Let us find the equation!
First, lets find the slope: point= (0,5) and (4,0)
m = (y(sub1) - y(sub2))/(x(sub1) - x(sub2))
m = (5-0)/(0-4)
m = (5)/(-4)
m = -5/4
Now, the equation: point=(0,5)
y - y(sub1) = m(x - x(sub1))
y - 5 = (-5/4)x
y = (-5/4)x + 5
This linear line doesn't cross the origin....
The origin is the place where you start from counting how much left on the x-axis or y-axis....
If you do not count any number up, down, left, or right, then you get the ordered point (0,0)